OpenStudy (anonymous):

solve for the variable. reduce the answer to the lowest terms 5h-3h+9h=176 h=???

OpenStudy (anonymous):

16

OpenStudy (anonymous):

\[11x = 176\] \[x = 16\]

OpenStudy (jiteshmeghwal9):

5h-3h+9h=176 2h+9h=176 11h=176 h=176/11=16

OpenStudy (anonymous):

Is Mahesh your brother?? @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):



OpenStudy (jiteshmeghwal9):

ya!@waterineyes

OpenStudy (anonymous):

That is nice...@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

Thanx!